Given: 𝑧 = 6 + 7𝑖

Calculate the following:

1) 𝑧-4

2) 𝑧1/3, graph the roots.

3) ln 𝑧 , for 𝑛 = 1, 3, 5 & 7

4) 𝑓(𝑧) = π‘§π‘§Μ…βˆ’ 𝑧̅𝑖 + 5𝑖, and check your answer using 𝑧 = π‘₯ + 𝑖y


1
Expert's answer
2021-02-16T05:47:28-0500

1) "z=6+7i\Rightarrow z^{-1}=\frac{1}{6+7i}=\frac{6-7i}{85}." Hence "z^{-4}=\frac{(6-7i)^4}{85^4}=\frac{-6827+i2184}{85^4}."

2) "z=\sqrt{85}(\frac{6}{\sqrt{85}}+i\frac{7}{\sqrt{85}})." Hence, "z=re^{i\theta}" where "r=\sqrt{85}, cos \ \theta=\frac{6}{\sqrt{85}}, sin \ \theta =\frac{7}{\sqrt{85}}." Hence "z^{1/3}=r^{1/3}e^{i\theta/3}, r^{1/3}e^{i(\theta/3+2\pi/3)}, r^{1/3}e^{i(\theta/3+4\pi/3)}." In other words, "z^{1/3}= \xi\ or\ \xi\omega \ or\ \xi \omega^2" where "\xi=r^{1/3}e^{i\theta/3} ." (Diagram is attached below) Here "\omega" denotes complex cube root of unity.


3) "ln (z)= ln |z|+i\theta+i2\pi n= ln \sqrt{85} +i\theta+i2\pi n." Now we put "n=1,3,5,7." to get, "ln (z)= ln \sqrt{85} +i\theta+i2\pi , ln \sqrt{85} +i\theta+i6\pi n, ln \sqrt{85} +i\theta+i10\pi n, ln \sqrt{85} +i\theta+i14\pi n."4) "f(z)=(6+7i)(6-7i)-(6-7i)i+5i=78-i."

Verification: "z\overline{z}=(x+iy)(x-iy)=x^2+y^2." Also "i\overline{z}=ix-y." Hence, "f(z)=x^2+y^2-ix-y+5i=x^2+y^2+y+i(5-x)." Hence "f(6+7i)=6^2+7^2-7+i(5-6)=" "78-i."